# The relevance of ‘on’ and ‘off’ transitions in quantum pair production experiments

### v1

@@ I Introduction
Particle creation phenomena are naturally described by quantum field theories in non-trivial backgrounds. Prominent examples include cosmological particle production resulting from the expansion of the Universe [1, 2, 3], Hawking radiation emitted by black holes [4, 5], and the Schwinger effect, which occurs in the presence of strong electromagnetic fields [6, 7, 8].
These scenarios are typically difficult to access experimentally. Instead, motivated by the original idea by Unruh [9], analog gravity experiments [10, 11, 12] have been used as a tool to explore the dynamics of quantum fields in non-trivial backgrounds or effective curved spacetimes. In recent years, numerous experiments have been carried out in hydrodynamical, condensed matter, optical systems, and others [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27], demonstrating the potential of such platforms for the study of quantum fields.
@@
Specifically, we study the impact of ‘on’ and ‘off’ transitions on particle production in $D$-dimensional Friedmann-Lemaître-Robertson-Walker (FLRW) expanding universes. We will consider a massive spectator scalar field, allowing for a non-minimal coupling to the geometry. Our results demonstrate that the coupling parameter between the field and the geometry has a strong influence on the resulting pair production spectrum. In particular, we find that particle production during the targeted time window is inevitably affected by ‘on’ and ‘off’ transition periods. Furthermore, in the case of a non-conformal coupling, production during abrupt transitions dominates pair creation, significantly overshadowing the contributions from the intermediate region. However, when the coupling is conformal, the ‘on’ and ‘off’ transitions do not substantially enhance pair creation, resulting in a much lower overall density of produced particles.
We will then apply these fundamental results to two experimental setups, highlighting the need for extreme caution when interpreting the physical results of these experiments. On the one hand, we will discuss analog pair production in (1+2)-dimensional Bose-Einstein condensates (BECs), which simulates the problem of a non-conformally coupled field in an FLRW universe. On the other hand, we will discuss the Schwinger effect due to a switchable electric field in $(1+3)$ dimensions. Its behavior regarding ‘on’ and ‘off’ transitions is for the most part equivalent to that of a conformally coupled field in an FLRW universe. However, unlike the cosmological case, the anisotropic nature of the electric field causes the contribution from the intermediate regime to increasingly dominate over that of the abrupt transitions when the field remains switched on for a sufficiently long duration, leading to enhanced particle production as the field duration grows.
@@ II Cosmological pair production
where $\eta$ is the conformal time and $a(\eta)$ the scale factor. The dynamics of a real, non-minimally coupled to gravity scalar field $\varphi(\eta,\bm{x})$ with mass $m$ is described by the equation
@@
The quantum vacuum is the state annihilated by all the annihilation operators, i.e., $\hat{b}_{\bm{k}}|0\rangle=0$ for all $\bm{k}$. The operators satisfy the standard commutation relations $[\hat{b}_{\bm{k}},\hat{b}_{\bm{k}^{\prime}}^{\dagger}]=\delta(\bm{k}-\bm{k}^{\prime})$, whereas all the other commutators vanish.
@@
The ‘in’ and ‘out’ bases are associated with the natural quantizations of observers living before and after the expansion of the universe. Due to the time-dependence of the geometry, even if the initial state of the system is the vacuum state as understood by the observer at $\eta\leq\eta_{\text{on}}$, it will be in general an excited state for another observer at $\eta\geq\eta_{\text{off}}$, where the particle content is non-vanishing. To measure this particle creation, we write the ‘out’ modes in terms of the ‘in’ modes,
@@
where $\alpha_{k}$ and $\beta_{k}$ are Bogoliubov coefficients [41, 42], which satisfy $|\alpha_{k}|^{2}-|\beta_{k}|^{2}=1$. The expectation number density of excited quanta per mode $k$ is precisely given by the square modulus of the $\beta$-coefficient,
@@
where this coefficient is time-independent and can be computed in terms of the ‘in’ and ‘out’ solutions:
@@
for $-\delta/2\leq\eta\leq\delta/2$, remaining constant outside this interval. $\Theta_{\delta}(\eta)$ is centered at the origin and its width is parametrized by $\delta$, such that for $\delta=0$ we recover the discontinuous Heaviside step function. We can then write the expansion rate during the entire expansion, including the transitions, as
@@
In Fig. 1, we represent this expansion rate for various transition durations $\delta$. In the following, we will say that the scale factor undergoes an abrupt transition when it is continuous but not differentiable. During these abrupt transitions, the expansion rate $a^{\prime}(\eta)$ involves discontinuous (but finite) step functions, as depicted in Fig. 1 for $\delta=0$.
@@
In Fig. 2, we calculate the total number density of produced pairs, as defined in Eq. (11), for fast transitions ($\delta=0.1/a_{0}^{\prime}$) and various durations of the intermediate expansion, $\eta_{\text{on}}\leq\eta\leq\eta_{\text{off}}$. Specifically, we fix $\eta_{\text{on}}=0$ and numerically compute the $\beta$-Bogoliubov coefficient using Eq. (10) for each process, with its duration parametrized by a particular value of $\eta_{\text{off}}$. We then integrate over all modes $k$ to obtain the total density. This procedure is repeated for each value of $\eta_{\text{off}}$. We examine how these particle densities depend on the coupling parameter $\xi$ and present results for $D=2$ and $D=3$ spatial dimensions. Our analysis reveals that 1) the total number density of produced pairs in the conformal coupling case, where $\xi=(D-1)/(4D)$, is significantly lower—by several orders of magnitude—than in the non-conformal case; and that 2) particle production stabilizes for sufficiently prolonged expansions.
@@
Figure 2: Total number density of produced particles $n$ as a function of $\eta_{\text{off}}$ ($\eta_{\text{on}}=0$) for fast transitions ($\delta=0.1/a_{0}^{\prime}$), in the case of two (left) and three (right) spatial dimensions. Results are shown for different values of the coupling $\xi$, where the red lines correspond to the conformal coupling case.
@@
Figure 3: Function $\mathcal{C}_{k}$ for $D=3$ dimensional universe expansion with $\Delta\eta=3/a_{0}$, for different transition durations $\delta$, fixed $k=a_{0}$. The left panel illustrates the conformal coupling case ($\xi=1/6$) while the right panel corresponds to a non-conformal coupling case ($\xi=1$).
@@
1) Production in the conformal case is suppressed with respect to the non-conformal case. The difference between these two cases can be understood by examining the behavior of the frequency (4), since pair production is dictated by the dynamics of the field modes. It is when the frequency rapidly varies over time, that particle production is enhanced. During the intermediate expansion phase ($\eta_{\text{on}}\leq\eta\leq\eta_{\text{off}}$), the frequency remains bounded in both cases, provided the expansion is sufficiently smooth. However, a crucial distinction arises during the abrupt transitions: while $a^{\prime}$ remains bounded, $a^{\prime\prime}$ approaches a Dirac delta. In the conformal coupling case, the mode frequency simplifies to $\omega_{k}^{2}=k^{2}+m^{2}a^{2}$, meaning that both $\omega_{k}$ and its time-derivative remain bounded even during abrupt transitions. Conversely, for non-conformal coupling, the frequency explicitly depends on the first and second derivatives of the scale factor. As a result, the frequency and its time-derivatives sharply diverge at the transition points, leading to significant enhancement of particle production. The crucial point is that this enhancement arises not from the intermediate expansion, but rather from the ‘on’ and ‘off’ transitions. This is particularly evident in our example, as in our computations we employ a constant expansion rate during the intermediate regime, ensuring that the derivatives of the scale factor entering the frequency (4) vanish except at the transitions [^1] Although the quantitative differences between conformal and non-conformal couplings may vary depending on the specific form of the scale factor, the qualitative behavior highlighted here is robust..
@@
2) Production stabilizes for sufficiently prolonged intermediate expansions. This behavior reinforces the conclusion that the dominant contribution to particle creation arises from the abrupt ‘on’ and ‘off’ transitions. Even when the intermediate expansion lasts significantly longer, the particle number density remains effectively constant, indicating that the intermediate phase contributes minimally to the overall production. The resulting horizontal asymptote can thus be interpreted as capturing the average effect of the abrupt transitions. We will see that, in the case of the Schwinger effect, anisotropies lead to a drastically different asymptotic behavior.
@@
×$\left.\times\left[\frac{a^{\prime\prime\prime}}{a}+(D-4)\frac{a^{\prime}a^{\prime\prime}}{a^{2}}-(D-3)\left(\frac{a^{\prime}}{a}\right)^{3}\right]\right|.$ (15)
@@
which is bounded from above by $\mathcal{C}_{k=0}=\mathcal{H}/m$. Even in cases where the scale factor undergoes abrupt transitions, the function $\mathcal{C}_{k}$ remains finite, as it depends only on the first derivative of the scale factor and not on higher orders. Nevertheless, higher-order time derivatives of the frequency involve higher derivatives of the scale factor, which, in the limit of abrupt transitions, tend to Dirac delta distributions and their derivatives. Consequently, although still relevant, the impact of transitions in pair production is smaller in this case than in the non-conformal scenario.
For the particular shape given in Eq. (13), Fig. 3 shows the function $\mathcal{C}_{k}$ for different values of $\delta$ in $D=3$ dimensions. In the conformal coupling case $\xi=1/6$, $\mathcal{C}_{k}$ remains bounded throughout the entire expansion, even during abrupt transitions. However, in the non-conformal coupling case ($\xi=1$), $\mathcal{C}_{k}$ exhibits sharp oscillations during the ‘on’ and ‘off’ transitions. The amplitude of these oscillations increases by several orders of magnitude as the transitions become shorter, highlighting the sensitivity of the system to more rapid transitions. It is clear in this case that the primary contribution to the particle excitation number arises predominantly from the transitions, overshadowing the effects of the linear expansion in the intermediate region.
@@
In Sec. IV and Sec. V we present two illustrative examples from laboratory settings where we apply the results just developed. The first involves gravitational analog experiments with BECs that mimic the dynamics of a non-conformally coupled scalar field in an FLRW expanding universe. The second focuses on the Schwinger effect, whose anisotropic nature introduces important nuances to our analysis that we discuss below.
@@ IV BEC analog experiment
We focus on the problem of analog particle production in a quasi two-dimensional, spin-0 BEC [48, 27, 28, 46]. Low-energy excitations on top of the condensate’s ground state behave as a massless scalar field propagating in a curved spacetime defined by the so-called acoustic metric. This acoustic metric, determined by the properties of the condensate, can be experimentally controlled to emulate a two-dimensional FLRW metric. As a result, the system provides an analog for cosmological particle production in a $(1+2)$-dimensional spacetime.
@@
This corresponds to the situation where the coupling is minimal and therefore non-conformal, and second derivatives of the scale factor appear in the frequency. Regarding the density of produced particles, this BEC experiment corresponds to the scenario described by the non-conformal coupling curve $\xi=0$ in the left panel of Fig. 2.
Figure 4: Function $\mathcal{C}_{k}$ for a typical analog gravity experiment in quasi two-dimensional BECs for $k=0.05\,\mu\text{m}^{-1}$ and transitions of different abruptness characterized by $\delta$. The values are normalized with respect to $\bar{\mathcal{C}}_{k}\simeq 5\times 10^{-8}$, assuming an intermediate region of duration $\Delta t=3\,\text{ms}$. The BEC parameters correspond to the expansion linear in $t$ presented in Ref. [27].
We computed the function $\mathcal{C}_{k}$ corresponding to the frequency (17), using the same functional form of the scale factor as in Eq. (13), but replacing conformal time $\eta$ with laboratory time $t$. In Fig. 4, we replicate the laboratory conditions reported in Ref. [27] for a scale factor linear in $t$ and observe that $\mathcal{C}_{k}$ increases by several orders of magnitude during the ‘on’ and ‘off’ transitions compared to its lowest value during the expansion, $\bar{\mathcal{C}}_{k}$. This reinforces our earlier conclusion: The effects of abrupt transitions overshadow the contributions from the intermediate expansion process, effectively masking the dynamics we aim to analyze. One must, therefore, be aware of the role of transitions concerning particle production when performing such experiments, as the outcome stems from the ‘on’ and ‘off’ transitions rather than from the background time-dependence in the intermediate region.

### v2

# Relevance of ‘on’ and ‘off’ transitions in quantum pair production experiments

@@ I Introduction
Particle creation phenomena are naturally described by quantum field theories in nontrivial backgrounds. Prominent examples include cosmological particle production resulting from the expansion of the Universe [1, 2, 3], Hawking radiation emitted by black holes [4, 5], and the Schwinger effect, which occurs in the presence of strong electromagnetic fields [6, 7, 8].
These scenarios are typically difficult to access experimentally. Instead, motivated by the original idea by Unruh [9], analog gravity experiments [10, 11, 12] have been used as a tool to explore the dynamics of quantum fields in nontrivial backgrounds or effective curved spacetimes. In recent years, numerous experiments have been carried out in hydrodynamical, condensed matter, optical systems, and others [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27], demonstrating the potential of such platforms for the study of quantum fields.
@@
Specifically, we study the impact of ‘on’ and ‘off’ transitions on particle production in $D$-dimensional Friedmann-Lemaître-Robertson-Walker (FLRW) expanding universes. We will consider a massive spectator scalar field, allowing for a nonminimal coupling to the geometry. Our results demonstrate that the coupling parameter between the field and the geometry has a strong influence on the resulting pair production spectrum. In particular, we find that particle production during the targeted time window is inevitably affected by ‘on’ and ‘off’ transition periods. Furthermore, in the case of a nonconformal coupling, production during abrupt transitions dominates pair creation, significantly overshadowing the contributions from the intermediate region. However, when the coupling is conformal, the ‘on’ and ‘off’ transitions do not substantially enhance pair creation, resulting in a much lower overall density of produced particles.
We will then apply these fundamental results to two experimental setups, highlighting the need for extreme caution when interpreting the physical results of these experiments. On the one hand, we will discuss analog pair production in (1+2)-dimensional Bose-Einstein condensates (BECs), which simulates the problem of a nonconformally coupled field in an FLRW universe. On the other hand, we will discuss the Schwinger effect due to a switchable electric field in $(1+3)$ dimensions. Its behavior regarding ‘on’ and ‘off’ transitions is for the most part equivalent to that of a conformally coupled field in an FLRW universe. However, unlike the cosmological case, the contribution from the intermediate regime increasingly dominates over that of the abrupt transitions when the field remains switched on for a sufficiently long time, leading to enhanced particle production as the field duration grows.
@@
We work in natural units, setting $c=\hbar=\varepsilon_{0}=1$.
@@ II Cosmological pair production
where $\eta$ is the conformal time and $a(\eta)$ the scale factor. The dynamics of a real, nonminimally coupled to gravity scalar field $\varphi(\eta,\bm{x})$ with mass $m$ is described by the equation
@@
The quantum vacuum is the state annihilated by all the annihilation operators, i.e., $\hat{b}_{\bm{k}}|0\rangle=0$ for all $\bm{k}$. The operators satisfy the standard commutation relations $\hat{b}_{\bm{k}},\hat{b}_{\bm{k}^{\prime}}^{\dagger}]=\delta(\bm{k}-\bm{k}^{\prime})$, whereas all the other commutators vanish.
@@
The ‘in’ and ‘out’ bases are associated with the natural quantizations of observers living before and after the expansion of the universe. Due to the time-dependence of the geometry, even if the initial state of the system is the vacuum state as understood by the observer at $\eta\leq\eta_{\text{on}}$, it will be in general an excited state for another observer at $\eta\geq\eta_{\text{off}}$, where the particle content is nonvanishing. To measure this particle creation, we write the ‘out’ modes in terms of the ‘in’ modes,
@@
where $\alpha_{k}$ and $\beta_{k}$ are Bogoliubov coefficients [41, 42], which satisfy $|\alpha_{k}|^{2}-|\beta_{k}|^{2}=1$. The expectation number density of excited quanta per mode $k$ is precisely given by the square modulus of the $\beta$ coefficient,
@@
where this coefficient is time independent and can be computed in terms of the ‘in’ and ‘out’ solutions,
@@
for $\delta/2\leq\eta\leq\delta/2$, remaining constant outside this interval. $\Theta_{\delta}(\eta)$ is centered at the origin and its width is parametrized by $\delta$, such that for $\delta=0$ we recover the discontinuous Heaviside step function. We can then write the expansion rate during the entire expansion, including the transitions, as
@@
In Fig. 1, we represent this expansion rate for various transition durations $\delta$. We restrict our analysis to transitions that are fast compared to the characteristic expansion rate in the intermediate region—for the specific scale factor in (13), this corresponds to $\delta a_{0}’\lesssim 1$. In the following, we will say that the scale factor undergoes an abrupt transition when it is continuous but not differentiable. During these abrupt transitions, the expansion rate $a^{\prime}(\eta)$ involves discontinuous (but finite) step functions, as depicted in Fig. 1 for $\delta=0$. In the subsequent figures, we fix the intermediate expansion rate to $a_{0}^{\prime}=m$.
@@
In Fig. 2, we calculate the total number density of produced pairs, as defined in Eq. (11), for various durations of the intermediate expansion, $\eta_{\text{on}}\leq\eta\leq\eta_{\text{off}}$, and for different transition durations $\delta$. Specifically, we fix $\eta_{\text{on}}=0$ and numerically compute the $\beta$-Bogoliubov coefficient using Eq. (10) for each process, with its duration parametrized by a particular value of $\eta_{\text{off}}$. We then integrate over all modes $k$ to obtain the total density. This procedure is repeated for each value of $\eta_{\text{off}}$. We examine how these particle densities depend on the coupling parameter $\xi$ and present results for $D=2$ and $D=3$ spatial dimensions. Our analysis reveals that 1) the total number density of produced pairs in the conformal coupling case, where $\xi=(D-1)/(4D)$, is significantly lower—by several orders of magnitude—than in the nonconformal case; 2) particle production stabilizes for sufficiently prolonged expansions; and 3) for nonconformal couplings, fast transitions yield a much larger asymptotic number density than slower transitions.
@@
Figure 2: Total number density of produced particles $n$ as a function of $\eta_{\text{off}}$ ($\eta_{\text{on}}=0$), in the case of two and three spatial dimensions. Results are shown for different values of the coupling $\xi$, where the red continuous lines correspond to the conformal coupling case. ‘On’ and ‘off’ transitions are chosen sufficiently fast ($\delta=0.1$) to guarantee convergence to the limiting behavior as the product $\delta a_{0}^{\prime}$ becomes small. We also consider slower transition rates, comparable to the expansion rate in the intermediate region ($\delta=1$). Here, $\delta$ and $\eta_{\text{off}}$ are expressed in units of $m^{-1}$, while the number density $n$ is given in units of $m^{3}$.
@@
Figure 3: Function $\mathcal{C}_{k}$ for $D=3$ dimensional universe expansion with $\Delta\eta=3$, for different transition durations $\delta$, fixed $k=1$. The left panel illustrates the conformal coupling case ($\xi=1/6$) while the right panel corresponds to a nonconformal coupling case ($\xi=1$). Here, $\eta$ and $k$ are expressed in units of $m^{-1}$ and $m$, respectively.
@@
1) Production in the conformal case is suppressed with respect to the nonconformal case. The difference between these two cases can be understood by examining the behavior of the frequency (4), since pair production is dictated by the dynamics of the field modes. It is when the frequency rapidly varies over time, that particle production is enhanced. During the intermediate expansion phase ($\eta_{\text{on}}\leq\eta\leq\eta_{\text{off}}$), the frequency remains bounded in both cases, provided the expansion is sufficiently smooth. However, a crucial distinction arises during the abrupt transitions: while $a^{\prime}$ remains bounded, $a^{\prime\prime}$ approaches a Dirac delta. In the conformal coupling case, the mode frequency simplifies to $\omega_{k}^{2}=k^{2}+m^{2}a^{2}$, meaning that both $\omega_{k}$ and its time derivative remain bounded even during abrupt transitions. Conversely, for nonconformal coupling, the frequency explicitly depends on the first and second derivatives of the scale factor. As a result, the frequency and its time derivatives sharply diverge at the transition points, leading to significant enhancement of particle production. The crucial point is that this enhancement arises not from the intermediate expansion, but rather from the ‘on’ and ‘off’ transitions. This is particularly evident in our example, as in our computations we employ a constant expansion rate during the intermediate regime, ensuring that the derivatives of the scale factor entering the frequency (4) vanish except at the transitions [^1] Although the quantitative differences between conformal and nonconformal couplings may vary depending on the specific form of the scale factor, the qualitative behavior highlighted here is robust..
@@
2) Production stabilizes for sufficiently prolonged intermediate expansions. Even when the intermediate expansion lasts significantly longer, the particle number density remains effectively constant, indicating that, from a certain point on, the intermediate phase contributes minimally to the overall production. We will see that the Schwinger effect exhibits a drastically different asymptotic behavior.
3) For nonconformal couplings, fast transitions yield a much larger asymptotic number density than slower transitions. This behavior reinforces the conclusion that the dominant contribution to particle creation arises from the abrupt ‘on’ and ‘off’ transitions. When the coupling is nonconformal, considering slower transitions in comparison to the expansion rate in the intermediate region significantly reduces production. The resulting horizontal asymptote in these cases can thus be interpreted as capturing mainly the average effect of the abrupt transitions for fast transitions. This effect is comparatively suppressed in the conformal coupling scenario, which tells us that transitions do not have a significant impact in this case, consistent with point 1).
@@
$\left.\times\left[\frac{a^{\prime\prime\prime}}{a}+(D-4)\frac{a^{\prime}a^{\prime\prime}}{a^{2}}-(D-3)\left(\frac{a^{\prime}}{a}\right)^{3}\right]\right|.$ (15)
@@
which is bounded from above by $\mathcal{C}_{k=0}=\mathcal{H}/m$. Even in cases where the scale factor undergoes abrupt transitions, the function $\mathcal{C}_{k}$ remains finite, as it depends only on the first derivative of the scale factor and not on higher orders. Nevertheless, higher-order time derivatives of the frequency involve higher derivatives of the scale factor, which, in the limit of abrupt transitions, tend to Dirac delta distributions and their derivatives. Consequently, although still relevant, the impact of transitions in pair production is smaller in this case than in the nonconformal scenario.
For the particular shape given in Eq. (13), Fig. 3 shows the function $\mathcal{C}_{k}$ for different values of $\delta$ in $D=3$ dimensions. In the conformal coupling case $\xi=1/6$, $\mathcal{C}_{k}$ remains bounded throughout the entire expansion, even during abrupt transitions. However, in the nonconformal coupling case ($\xi=1$), $\mathcal{C}_{k}$ exhibits sharp oscillations during the ‘on’ and ‘off’ transitions. The amplitude of these oscillations increases by several orders of magnitude as the transitions become shorter, highlighting the sensitivity of the system to more rapid transitions. It is clear in this case that the primary contribution to the particle excitation number arises predominantly from the transitions, overshadowing the effects of the linear expansion in the intermediate region.
@@
In Sec. IV and Sec. V we present two illustrative examples from laboratory settings where we apply the results just developed. The first involves gravitational analog experiments with BECs that mimic the dynamics of a nonconformally coupled scalar field in an FLRW expanding universe. The second focuses on the Schwinger effect, whose anisotropic nature introduces important nuances to our analysis that we discuss below.
@@ IV BEC analog experiment
We focus on the problem of analog particle production in a quasi-two-dimensional, spin-0 BEC [48, 27, 28, 46]. Low-energy excitations on top of the condensate’s ground state behave as a massless scalar field propagating in a curved spacetime defined by the so-called acoustic metric. This acoustic metric, determined by the properties of the condensate, can be experimentally controlled to emulate a two-dimensional FLRW metric. As a result, the system provides an analog for cosmological particle production in a $(1+2)$-dimensional spacetime.
@@
This corresponds to the situation where the coupling is minimal and therefore nonconformal, and second derivatives of the scale factor appear in the frequency. Regarding the density of produced particles, this BEC experiment corresponds to the scenario described by the non-conformal coupling curve $\xi=0$ in the leftmost panel of Fig. 2.
Figure 4: Function $\mathcal{C}_{k}$ for a typical analog gravity experiment in quasi-two-dimensional BECs for $k=0.5\,\mu\text{m}^{-1}$ and transitions of different abruptness characterized by $\delta$. The values are normalized with respect to $\bar{\mathcal{C}}_{k}\simeq 5\times 10^{-11}$, assuming an intermediate region of duration $\Delta t=3\,\text{ms}$. The BEC parameters correspond to the expansion linear in $t$ presented in Ref. [27].
We computed the function $\mathcal{C}_{k}$ corresponding to the frequency (17), using the same functional form of the scale factor as in Eq. (13), but replacing conformal time $\eta$ with laboratory time $t$. In Fig. 4, we replicate the laboratory conditions reported in Ref. [27] for a scale factor linear in $t$ and observe that $\mathcal{C}_{k}$ increases by several orders of magnitude during the ‘on’ and ‘off’ transitions compared to its lowest value during the expansion, $\bar{\mathcal{C}}_{k}$. This reinforces our earlier conclusion: The effects of abrupt transitions overshadow the contributions from the intermediate expansion process, effectively masking the dynamics we aim to analyze. One must, therefore, be aware of the role of transitions concerning particle production when performing such experiments, as the number of particles created stems from the ‘on’ and ‘off’ transitions rather than from the background time-dependence in the intermediate region.